The product of 3 and the sum of a number and 5 is the same as the product of -2 and the difference of the same number from 3. Find the number.
Answer by ankor@dixie-net.com(22740)

The product of 3 and the sum of a number and 5 is the same as the product of -2 and the difference of the same number from 3. Find the number.
:
Let x = " a number"
:
An equation for the statement:
"The product of 3 and the sum of a number and 5 is the same as the product of -2 and the difference of the same number from 3."
3(x+5) = -2(3-x)
;
3x + 15 = -6 + 2x
:
3x - 2x = -6 -15
:
x = -21
:
:
Try this in the equation:
3(-21+5) = -2(3 - (-21)); minus a minus is a plus
3(-16) = -2(24)
-48 = -48